# Mitigating Label Noise using Prompt-Based Hyperbolic Meta-Learning in Open-Set Domain Generalization

### v1

@@ 5.1 Implementation Details
TCL [26] 58.32 59.21 51.72 53.66 48.28 42.91 46.78 38.29 31.74 31.55 22.88 24.30 47.58 42.17 37.67
NPN [55] 64.30 70.87 61.99 51.66 52.10 45.40 55.65 44.88 38.64 35.58 22.35 25.86 51.80 47.55 42.97
BadLabel [75] 54.93 55.73 48.24 53.72 53.25 46.55 50.23 55.36 45.70 31.55 21.84 28.38 47.61 46.55 42.22
DISC [43] 53.47 56.13 47.22 54.47 47.46 43.48 53.27 53.97 44.33 24.01 16.75 11.52 46.31 43.58 36.64
LSL [29] 58.97 58.93 52.15 49.97 48.17 39.20 47.50 44.07 34.63 30.59 12.44 16.81 46.76 40.90 35.70
PLM [78] 55.57 42.33 38.27 41.78 43.09 32.95 45.75 40.44 33.26 33.27 12.11 15.46 38.33 34.49 29.99
ARPL [3] 62.52 67.96 59.46 52.35 45.29 41.09 50.13 44.47 37.01 29.56 13.49 22.70 48.64 42.80 40.07
ODGNet [4] 63.00 70.61 61.18 58.08 40.01 44.97 58.33 53.37 48.89 22.84 9.69 16.48 50.56 43.42 42.88
MLDG [56] 60.26 69.11 59.35 58.66 55.83 49.03 58.07 51.18 45.08 25.87 18.48 16.40 37.22 48.65 42.47
SWAD [6] 59.94 69.23 58.69 49.59 48.04 40.04 37.44 34.32 25.96 19.10 20.72 12.86 41.52 41.52 34.39
MixStyle [83] 60.10 65.39 56.89 55.16 44.70 44.01 59.31 47.35 39.93 34.54 17.49 20.86 52.28 43.73 40.42
MEDIC-cls [66] 62.20 52.63 53.23 54.60 54.05 46.51 59.31 52.02 47.65 34.54 28.22 21.44 52.66 46.73 41.96
MEDIC-bcls [66] 62.20 57.47 53.93 54.60 53.10 46.38 59.31 53.70 48.68 34.54 32.71 24.06 52.66 49.25 42.76
@@
Table 2: Results (%) of PACS on ResNet18 [24]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $50\%$ is selected.
@@
TCL [26] 54.68 52.40 46.51 52.78 22.65 30.94 47.19 37.73 34.89 26.33 9.83 7.62 45.25 30.65 29.99
NPN [55] 48.38 38.12 33.55 35.71 32.33 24.47 38.94 26.88 18.60 26.93 27.59 18.96 37.49 31.23 23.90
BadLabel [75] 46.20 57.45 45.07 45.34 47.29 37.89 35.17 43.35 32.14 28.40 26.95 15.71 38.78 43.76 32.70
DISC [43] 52.52 56.07 50.55 46.84 31.91 30.35 28.47 28.28 19.97 30.83 25.63 24.78 39.67 35.47 31.41
LSL [29] 41.36 30.83 20.27 42.28 39.78 31.40 42.39 37.59 30.89 26.90 15.40 7.42 38.23 30.90 22.50
PLM [78] 55.57 42.33 38.27 39.21 27.81 24.93 33.01 27.81 21.49 25.52 6.65 13.41 38.33 26.15 24.53
ARPL [3] 55.41 62.40 54.17 45.72 44.50 34.51 43.73 38.44 30.13 27.30 7.65 20.95 43.04 38.25 34.94
ODGNet [4] 60.66 63.57 56.75 55.09 40.01 44.97 46.52 39.85 32.10 32.02 24.40 17.09 48.57 41.96 37.73
MLDG [56] 59.37 68.02 58.54 56.49 50.15 44.92 46.78 46.02 36.91 23.69 24.32 16.40 46.58 47.13 39.19
SWAD [6] 58.58 67.77 56.25 45.78 41.39 38.30 34.19 33.89 23.95 20.43 14.15 6.81 39.75 39.30 31.33
MixStyle [83] 54.04 62.25 30.23 41.78 37.68 27.03 47.09 26.67 27.03 30.88 22.81 17.09 43.45 37.35 25.35
MEDIC-cls [66] 60.58 51.37 44.29 53.28 51.88 44.12 50.54 49.07 42.84 36.67 28.00 20.83 50.27 45.08 38.02
MEDIC-bcls [66] 60.58 48.99 43.25 53.28 37.32 33.99 50.54 44.08 39.39 36.67 30.58 21.83 50.27 40.24 34.62
@@
Table 3: Results (%) of PACS on ResNet18 [24]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $80\%$ is selected.
@@
TCL [26] 31.58 25.81 17.39 27.08 26.60 16.45 27.69 27.17 17.66 21.20 8.52 14.87 26.89 22.03 16.59
NPN [55] 17.21 12.49 10.18 24.27 12.67 13.87 22.85 12.99 10.85 19.66 4.31 11.63 21.00 10.62 11.63
BadLabel [75] 22.62 14.11 22.62 15.95 14.50 10.16 19.39 24.34 14.95 26.13 14.14 17.96 21.02 16.77 16.42
DISC [43] 22.05 19.53 15.27 24.77 23.65 17.53 27.13 22.49 14.76 16.03 12.86 10.00 22.05 19.63 14.39
LSL [29] 18.58 22.82 13.89 23.64 16.71 14.56 15.37 15.84 7.94 21.68 1.92 8.26 19.82 14.32 11.16
PLM [78] 24.39 8.71 9.41 30.08 24.67 17.85 20.94 13.16 12.21 21.76 21.26 11.74 24.29 16.95 12.80
ARPL [3] 38.77 23.79 15.88 22.12 20.58 11.40 23.98 14.45 8.98 25.76 16.45 11.71 27.66 18.82 11.99
ODGNet [4] 31.18 19.56 18.49 27.64 6.64 12.81 20.78 21.43 12.81 21.65 22.00 7.9 25.31 17.41 13.00
MLDG [56] 33.04 9.18 12.11 22.45 19.28 11.18 28.16 23.38 13.64 23.19 4.47 16.40 26.71 14.08 13.33
SWAD [6] 18.09 18.69 10.19 22.51 20.22 11.97 23.67 21.96 11.89 19.75 12.20 15.29 21.01 18.27 12.33
MixStyle [83] 25.28 22.05 16.88 24.70 17.68 12.90 21.61 20.39 11.14 24.44 12.39 15.13 24.01 18.13 14.01
MEDIC-cls [66] 30.61 15.03 21.20 22.33 22.47 14.20 29.55 26.02 14.96 23.11 15.61 8.74 26.40 19.78 14.78
MEDIC-bcls [66] 30.61 12.82 11.92 22.33 21.15 11.45 29.55 22.67 13.82 23.11 8.34 7.47 26.40 16.25 11.17
@@
Table 4: Results (%) of PACS on ResNet18 [24]. The open-set ratio is $6{:}1$ and asymmetric label noise with ratio $50\%$ is selected.
@@
TCL [26] 15.83 3.33 10.69 35.21 30.54 20.96 26.46 24.24 14.88 20.35 1.2 2.63 24.46 14.83 12.29
NPN [55] 44.43 37.82 28.94 38.65 39.48 23.11 31.92 20.12 13.49 23.24 11.25 4.29 34.56 27.17 17.46
BadLabel [75] 37.16 44.56 35.70 26.89 31.98 24.77 31.10 34.52 25.33 13.88 16.21 20.38 27.26 31.82 26.54
DISC [43] 46.20 42.31 38.82 44.28 42.69 34.41 42.70 35.22 26.28 32.77 2.77 22.37 41.49 30.75 30.47
LSL [29] 27.30 19.63 14.32 24.27 11.56 13.18 25.63 14.77 12.04 22.21 16.31 25.07 24.85 15.57 16.15
PLM [78] 17.37 23.20 14.13 25.45 24.44 15.04 20.06 9.68 11.93 20.48 2.60 3.24 20.84 14.98 11.09
ARPL [3] 38.69 28.90 31.88 37.71 29.82 19.55 33.99 24.57 16.20 20.56 26.92 17.79 32.74 27.55 21.35
ODGNet [4] 45.15 49.11 39.31 37.59 34.82 24.43 42.96 42.10 24.43 26.00 15.02 15.65 37.93 35.26 25.96
MLDG [56] 51.21 40.25 45.11 42.21 32.24 26.34 44.46 38.66 30.80 24.36 1.32 15.99 40.56 28.12 29.56
SWAD [6] 40.47 14.21 35.14 32.15 18.26 9.48 20.06 12.52 8.53 20.48 11.83 5.47 28.29 14.20 14.66
MixStyle [83] 49.76 41.10 40.39 35.96 36.32 26.54 41.72 32.04 23.55 25.58 4.81 23.55 38.25 28.57 28.51
MEDIC-cls [66] 46.20 45.01 37.13 37.46 29.69 22.92 36.41 30.80 20.83 31.07 23.88 13.29 37.79 32.35 23.54
MEDIC-bcls [66] 46.20 48.97 39.86 37.46 20.65 17.95 36.41 32.06 22.69 31.07 25.23 15.23 37.79 31.73 23.93
@@
Table 5: Results (%) of PACS on ViT-Base [16]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $20\%$ is selected.
@@
TCL [26] 67.77 79.17 67.64 64.35 63.45 56.91 51.32 50.24 42.30 26.43 28.25 20.54 52.47 55.28 46.85
NPN [55] 63.97 71.61 62.53 57.22 51.43 45.95 47.65 40.11 32.85 21.86 13.01 12.71 47.68 44.04 38.51
BadLabel [75] 62.66 64.81 57.87 51.59 60.49 49.13 50.85 58.42 47.75 31.15 18.54 29.40 49.06 50.57 46.04
DISC [43] 66.69 71.98 65.67 54.10 29.80 41.07 53.48 41.72 40.83 34.57 16.94 24.16 52.21 40.11 42.93
LSL [29] 67.04 71.46 63.82 63.10 63.42 58.01 53.27 54.92 47.14 28.44 4.37 27.51 52.96 48.54 49.12
PLM [78] 65.73 61.01 59.76 65.73 61.01 59.76 52.71 48.17 46.32 27.57 17.94 20.62 52.94 47.03 46.62
ARPL [3] 68.09 75.04 67.31 56.91 51.96 44.50 59.93 60.98 54.02 36.46 7.36 27.55 55.35 48.84 48.35
ODGNet [4] 66.48 71.17 65.60 62.44 65.75 58.60 60.50 59.99 53.54 30.14 4.05 15.68 54.89 50.24 48.36
MLDG [56] 66.16 72.74 64.05 57.85 54.34 47.79 60.80 61.93 54.69 36.38 11.21 29.16 55.30 50.06 48.92
SWAD [6] 63.00 72.01 62.08 64.79 66.60 60.22 56.68 59.18 51.10 29.90 20.51 22.51 53.59 54.58 48.98
MixStyle [83] 68.42 63.75 59.88 63.23 61.42 56.49 51.99 54.34 45.19 28.36 3.13 12.79 53.00 45.66 43.59
MEDIC-cls [66] 65.83 70.15 62.11 66.04 55.36 57.33 56.42 59.05 51.07 38.76 26.01 21.45 56.76 52.64 47.99
MEDIC-bcls [66] 65.83 70.08 63.40 66.04 55.18 55.40 56.42 54.09 49.89 38.76 14.81 25.12 56.76 48.54 48.45
@@
Table 6: Results (%) of PACS on ViT-Base [16]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $50\%$ is selected.
@@
TCL [26] 68.17 62.70 67.56 64.48 67.84 61.30 45.80 35.84 35.68 22.29 9.28 3.54 50.19 43.92 42.02
NPN [55] 27.71 10.79 8.48 39.34 33.88 32.40 40.39 44.98 34.61 20.80 26.62 19.84 32.06 29.07 23.83
BadLabel [75] 51.62 61.25 50.41 32.52 37.26 26.67 43.58 55.85 42.45 31.60 28.24 26.93 39.83 45.65 36.62
DISC [43] 57.35 52.08 44.73 38.02 37.36 28.09 30.84 31.69 22.68 20.72 16.32 16.20 36.73 34.36 27.93
LSL [29] 65.99 68.68 62.18 55.66 49.85 44.97 49.20 44.94 36.86 29.90 6.91 19.18 50.19 42.60 40.80
PLM [78] 57.67 50.81 51.87 51.41 46.57 41.82 39.66 39.47 31.89 19.95 12.93 15.75 42.17 37.45 35.33
ARPL [3] 57.27 62.95 53.89 39.21 37.92 29.78 51.62 53.40 45.17 33.03 2.77 17.23 45.28 39.26 36.52
ODGNet [4] 68.09 76.73 67.32 64.79 64.09 59.64 53.22 52.74 47.68 34.39 20.16 18.61 55.12 53.43 48.31
MLDG [56] 67.29 75.04 65.97 66.98 66.41 62.00 55.44 54.65 47.35 33.83 15.82 23.25 55.89 52.98 49.64
SWAD [6] 68.58 78.86 68.33 63.29 65.39 58.28 51.68 52.20 44.65 26.48 24.19 16.84 52.51 55.16 47.03
MixStyle [83] 54.04 53.94 46.23 52.72 46.49 39.10 37.65 29.78 22.47 20.38 2.35 13.29 41.20 33.14 30.27
MEDIC-cls [66] 62.76 69.29 60.03 63.10 62.25 55.06 56.42 56.30 49.08 32.16 8.13 24.16 53.61 48.99 47.08
MEDIC-bcls [66] 62.76 50.05 44.19 63.10 45.67 46.56 56.42 39.96 40.02 32.16 27.74 23.91 53.61 40.86 38.67
@@
Table 7: Results (%) of PACS on ViT-Base [16]. The open-set ratio is $6{:}1$ and symmetric label noise with ratio $80\%$ is selected.

### v2

@@ 5.2 Evaluation Metrics
TCL [huang2023twin] 58.32 59.21 51.72 53.66 48.28 42.91 46.78 38.29 31.74 31.55 22.88 24.30 47.58 42.17 37.67
NPN [sheng2024adaptive] 64.30 70.87 61.99 51.66 52.10 45.40 55.65 44.88 38.64 35.58 22.35 25.86 51.80 47.55 42.97
BadLabel [zhang2024badlabel] 54.93 55.73 48.24 53.72 53.25 46.55 50.23 55.36 45.70 31.55 21.84 28.38 47.61 46.55 42.22
DISC [li2023disc] 53.47 56.13 47.22 54.47 47.46 43.48 53.27 53.97 44.33 24.01 16.75 11.52 46.31 43.58 36.64
LSL [kim2024learning] 58.97 58.93 52.15 49.97 48.17 39.20 47.50 44.07 34.63 30.59 12.44 16.81 46.76 40.90 35.70
PLM [zhao2024estimating] 55.57 42.33 38.27 41.78 43.09 32.95 45.75 40.44 33.26 33.27 12.11 15.46 38.33 34.49 29.99
ARPL [bendale2016towards] 62.52 67.96 59.46 52.35 45.29 41.09 50.13 44.47 37.01 29.56 13.49 22.70 48.64 42.80 40.07
ODGNet [bose2023beyond] 63.00 70.61 61.18 58.08 40.01 44.97 58.33 53.37 48.89 22.84 9.69 16.48 50.56 43.42 42.88
MLDG [shu2019meta] 60.26 69.11 59.35 58.66 55.83 49.03 58.07 51.18 45.08 25.87 18.48 16.40 37.22 48.65 42.47
SWAD [cha2021swad] 59.94 69.23 58.69 49.59 48.04 40.04 37.44 34.32 25.96 19.10 20.72 12.86 41.52 41.52 34.39
MixStyle [zhou2020domain] 60.10 65.39 56.89 55.16 44.70 44.01 59.31 47.35 39.93 34.54 17.49 20.86 52.28 43.73 40.42
MEDIC-cls [wang2023generalizable] 62.20 52.63 53.23 54.60 54.05 46.51 59.31 52.02 47.65 34.54 28.22 21.44 52.66 46.73 41.96
MEDIC-bcls [wang2023generalizable] 62.20 57.47 53.93 54.60 53.10 46.38 59.31 53.70 48.68 34.54 32.71 24.06 52.66 49.25 42.76
EBiL-HaDS-cls [peng2024advancing] 65.19 58.09 57.84 53.28 47.07 40.36 57.56 52.17 45.95 37.52 28.83 22.31 53.39 46.54 41.62
EBiL-HaDS-bcls [peng2024advancing] 65.19 63.82 60.63 53.28 46.70 39.80 57.56 50.58 45.63 37.52 30.61 26.55 53.39 47.93 43.15
@@
Table 2: Results (%) of PACS on ResNet18 [he2016deep]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $50\%$ is selected.
@@
TCL [huang2023twin] 54.68 52.40 46.51 52.78 22.65 30.94 47.19 37.73 34.89 26.33 9.83 7.62 45.25 30.65 29.99
NPN [sheng2024adaptive] 48.38 38.12 33.55 35.71 32.33 24.47 38.94 26.88 18.60 26.93 27.59 18.96 37.49 31.23 23.90
BadLabel [zhang2024badlabel] 46.20 57.45 45.07 45.34 47.29 37.89 35.17 43.35 32.14 28.40 26.95 15.71 38.78 43.76 32.70
DISC [li2023disc] 52.52 56.07 50.55 46.84 31.91 30.35 28.47 28.28 19.97 30.83 25.63 24.78 39.67 35.47 31.41
LSL [kim2024learning] 41.36 30.83 20.27 42.28 39.78 31.40 42.39 37.59 30.89 26.90 15.40 7.42 38.23 30.90 22.50
PLM [zhao2024estimating] 55.57 42.33 38.27 39.21 27.81 24.93 33.01 27.81 21.49 25.52 6.65 13.41 38.33 26.15 24.53
ARPL [bendale2016towards] 55.41 62.40 54.17 45.72 44.50 34.51 43.73 38.44 30.13 27.30 7.65 20.95 43.04 38.25 34.94
ODGNet [bose2023beyond] 60.66 63.57 56.75 55.09 40.01 44.97 46.52 39.85 32.10 32.02 24.40 17.09 48.57 41.96 37.73
MLDG [shu2019meta] 59.37 68.02 58.54 56.49 50.15 44.92 46.78 46.02 36.91 23.69 24.32 16.40 46.58 47.13 39.19
SWAD [cha2021swad] 58.58 67.77 56.25 45.78 41.39 38.30 34.19 33.89 23.95 20.43 14.15 6.81 39.75 39.30 31.33
MixStyle [zhou2020domain] 54.04 62.25 30.23 41.78 37.68 27.03 47.09 26.67 27.03 30.88 22.81 17.09 43.45 37.35 25.35
MEDIC-cls [wang2023generalizable] 60.58 51.37 44.29 53.28 51.88 44.12 50.54 49.07 42.84 36.67 28.00 20.83 50.27 45.08 38.02
MEDIC-bcls [wang2023generalizable] 60.58 48.99 43.25 53.28 37.32 33.99 50.54 44.08 39.39 36.67 30.58 21.83 50.27 40.24 34.62
EBiL-HaDS-cls [peng2024advancing] 61.15 62.20 54.97 52.47 43.90 37.71 49.66 48.05 40.75 34.39 28.62 20.98 49.42 45.69 38.60
EBiL-HaDS-bcls [peng2024advancing] 61.15 25.32 48.79 52.47 42.61 36.20 49.66 49.13 41.34 34.39 21.33 21.70 49.42 34.60 37.00
@@
Table 3: Results (%) of PACS on ResNet18 [he2016deep]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $80\%$ is selected.
@@
TCL [huang2023twin] 31.58 25.81 17.39 27.08 26.60 16.45 27.69 27.17 17.66 21.20 8.52 14.87 26.89 22.03 16.59
NPN [sheng2024adaptive] 17.21 12.49 10.18 24.27 12.67 13.87 22.85 12.99 10.85 19.66 4.31 11.63 21.00 10.62 11.63
BadLabel [zhang2024badlabel] 22.62 14.11 22.62 15.95 14.50 10.16 19.39 24.34 14.95 26.13 14.14 17.96 21.02 16.77 16.42
DISC [li2023disc] 22.05 19.53 15.27 24.77 23.65 17.53 27.13 22.49 14.76 16.03 12.86 10.00 22.05 19.63 14.39
LSL [kim2024learning] 18.58 22.82 13.89 23.64 16.71 14.56 15.37 15.84 7.94 21.68 1.92 8.26 19.82 14.32 11.16
PLM [zhao2024estimating] 24.39 8.71 9.41 30.08 24.67 17.85 20.94 13.16 12.21 21.76 21.26 11.74 24.29 16.95 12.80
ARPL [bendale2016towards] 38.77 23.79 15.88 22.12 20.58 11.40 23.98 14.45 8.98 25.76 16.45 11.71 27.66 18.82 11.99
ODGNet [bose2023beyond] 31.18 19.56 18.49 27.64 6.64 12.81 20.78 21.43 12.81 21.65 22.00 7.9 25.31 17.41 13.00
MLDG [shu2019meta] 33.04 9.18 12.11 22.45 19.28 11.18 28.16 23.38 13.64 23.19 4.47 16.40 26.71 14.08 13.33
SWAD [cha2021swad] 18.09 18.69 10.19 22.51 20.22 11.97 23.67 21.96 11.89 19.75 12.20 15.29 21.01 18.27 12.33
MixStyle [zhou2020domain] 25.28 22.05 16.88 24.70 17.68 12.90 21.61 20.39 11.14 24.44 12.39 15.13 24.01 18.13 14.01
MEDIC-cls [wang2023generalizable] 30.61 15.03 21.20 22.33 22.47 14.20 29.55 26.02 14.96 23.11 15.61 8.74 26.40 19.78 14.78
MEDIC-bcls [wang2023generalizable] 30.61 12.82 11.92 22.33 21.15 11.45 29.55 22.67 13.82 23.11 8.34 7.47 26.40 16.25 11.17
EBiL-HaDS-cls [peng2024advancing] 40.06 39.36 34.58 19.51 3.89 5.30 29.40 26.25 18.44 25.44 23.73 18.01 28.60 23.31 19.08
EBiL-HaDS-bcls [peng2024advancing] 40.06 15.06 23.84 19.51 12.30 11.93 29.40 29.83 18.73 25.44 26.37 17.71 28.60 20.89 18.05
@@
Table 4: Results (%) of PACS on ResNet18 [he2016deep]. The open-set ratio is 6:16{:}1 and asymmetric label noise with a ratio of $50\%$ is selected.
@@
TCL [huang2023twin] 15.83 3.33 10.69 35.21 30.54 20.96 26.46 24.24 14.88 20.35 1.2 2.63 24.46 14.83 12.29
NPN [sheng2024adaptive] 44.43 37.82 28.94 38.65 39.48 23.11 31.92 20.12 13.49 23.24 11.25 4.29 34.56 27.17 17.46
BadLabel [zhang2024badlabel] 37.16 44.56 35.70 26.89 31.98 24.77 31.10 34.52 25.33 13.88 16.21 20.38 27.26 31.82 26.54
DISC [li2023disc] 46.20 42.31 38.82 44.28 42.69 34.41 42.70 35.22 26.28 32.77 2.77 22.37 41.49 30.75 30.47
LSL [kim2024learning] 27.30 19.63 14.32 24.27 11.56 13.18 25.63 14.77 12.04 22.21 16.31 25.07 24.85 15.57 16.15
PLM [zhao2024estimating] 17.37 23.20 14.13 25.45 24.44 15.04 20.06 9.68 11.93 20.48 2.60 3.24 20.84 14.98 11.09
ARPL [bendale2016towards] 38.69 28.90 31.88 37.71 29.82 19.55 33.99 24.57 16.20 20.56 26.92 17.79 32.74 27.55 21.35
ODGNet [bose2023beyond] 45.15 49.11 39.31 37.59 34.82 24.43 42.96 42.10 24.43 26.00 15.02 15.65 37.93 35.26 25.96
MLDG [shu2019meta] 51.21 40.25 45.11 42.21 32.24 26.34 44.46 38.66 30.80 24.36 1.32 15.99 40.56 28.12 29.56
SWAD [cha2021swad] 40.47 14.21 35.14 32.15 18.26 9.48 20.06 12.52 8.53 20.48 11.83 5.47 28.29 14.20 14.66
MixStyle [zhou2020domain] 49.76 41.10 40.39 35.96 36.32 26.54 41.72 32.04 23.55 25.58 4.81 23.55 38.25 28.57 28.51
MEDIC-cls [wang2023generalizable] 46.20 45.01 37.13 37.46 29.69 22.92 36.41 30.80 20.83 31.07 23.88 13.29 37.79 32.35 23.54
MEDIC-bcls [wang2023generalizable] 46.20 48.97 39.86 37.46 20.65 17.95 36.41 32.06 22.69 31.07 25.23 15.23 37.79 31.73 23.93
EBiL-HaDS-cls [peng2024advancing] 54.93 56.65 48.44 31.58 30.11 21.19 37.13 32.23 22.83 25.44 23.73 18.01 37.27 35.68 27.62
EBiL-HaDS-bcls [peng2024advancing] 54.93 30.36 40.79 31.58 30.97 18.73 37.13 32.56 23.21 25.44 26.37 17.71 37.27 30.07 25.11
@@
Table 5: Results (%) of PACS on ViT-Base [dosovitskiy2021an]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $20\%$ is selected.
@@
TCL [huang2023twin] 67.77 79.17 67.64 64.35 63.45 56.91 51.32 50.24 42.30 26.43 28.25 20.54 52.47 55.28 46.85
NPN [sheng2024adaptive] 63.97 71.61 62.53 57.22 51.43 45.95 47.65 40.11 32.85 21.86 13.01 12.71 47.68 44.04 38.51
BadLabel [zhang2024badlabel] 62.66 64.81 57.87 51.59 60.49 49.13 50.85 58.42 47.75 31.15 18.54 29.40 49.06 50.57 46.04
DISC [li2023disc] 66.69 71.98 65.67 54.10 29.80 41.07 53.48 41.72 40.83 34.57 16.94 24.16 52.21 40.11 42.93
LSL [kim2024learning] 67.04 71.46 63.82 63.10 63.42 58.01 53.27 54.92 47.14 28.44 4.37 27.51 52.96 48.54 49.12
PLM [zhao2024estimating] 65.73 61.01 59.76 65.73 61.01 59.76 52.71 48.17 46.32 27.57 17.94 20.62 52.94 47.03 46.62
ARPL [bendale2016towards] 68.09 75.04 67.31 56.91 51.96 44.50 59.93 60.98 54.02 36.46 7.36 27.55 55.35 48.84 48.35
ODGNet [bose2023beyond] 66.48 71.17 65.60 62.44 65.75 58.60 60.50 59.99 53.54 30.14 4.05 15.68 54.89 50.24 48.36
MLDG [shu2019meta] 66.16 72.74 64.05 57.85 54.34 47.79 60.80 61.93 54.69 36.38 11.21 29.16 55.30 50.06 48.92
SWAD [cha2021swad] 63.00 72.01 62.08 64.79 66.60 60.22 56.68 59.18 51.10 29.90 20.51 22.51 53.59 54.58 48.98
MixStyle [zhou2020domain] 68.42 63.75 59.88 63.23 61.42 56.49 51.99 54.34 45.19 28.36 3.13 12.79 53.00 45.66 43.59
MEDIC-cls [wang2023generalizable] 65.83 70.15 62.11 66.04 55.36 57.33 56.42 59.05 51.07 38.76 26.01 21.45 56.76 52.64 47.99
MEDIC-bcls [wang2023generalizable] 65.83 70.08 63.40 66.04 55.18 55.40 56.42 54.09 49.89 38.76 14.81 25.12 56.76 48.54 48.45
EBiL-HaDS-cls [peng2024advancing] 65.43 55.59 50.68 65.48 65.80 60.71 56.83 58.38 50.43 37.22 18.15 24.87 56.24 49.48 46.67
EBiL-HaDS-bcls [peng2024advancing] 65.43 53.39 55.16 65.48 59.68 58.33 56.83 56.59 48.76 37.22 19.38 23.75 56.24 47.26 46.50
@@
Table 6: Results (%) of PACS on ViT-Base [dosovitskiy2021an]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $50\%$ is selected.
@@
TCL [huang2023twin] 68.17 62.70 67.56 64.48 67.84 61.30 45.80 35.84 35.68 22.29 9.28 3.54 50.19 43.92 42.02
NPN [sheng2024adaptive] 27.71 10.79 8.48 39.34 33.88 32.40 40.39 44.98 34.61 20.80 26.62 19.84 32.06 29.07 23.83
BadLabel [zhang2024badlabel] 51.62 61.25 50.41 32.52 37.26 26.67 43.58 55.85 42.45 31.60 28.24 26.93 39.83 45.65 36.62
DISC [li2023disc] 57.35 52.08 44.73 38.02 37.36 28.09 30.84 31.69 22.68 20.72 16.32 16.20 36.73 34.36 27.93
LSL [kim2024learning] 65.99 68.68 62.18 55.66 49.85 44.97 49.20 44.94 36.86 29.90 6.91 19.18 50.19 42.60 40.80
PLM [zhao2024estimating] 57.67 50.81 51.87 51.41 46.57 41.82 39.66 39.47 31.89 19.95 12.93 15.75 42.17 37.45 35.33
ARPL [bendale2016towards] 57.27 62.95 53.89 39.21 37.92 29.78 51.62 53.40 45.17 33.03 2.77 17.23 45.28 39.26 36.52
ODGNet [bose2023beyond] 68.09 76.73 67.32 64.79 64.09 59.64 53.22 52.74 47.68 34.39 20.16 18.61 55.12 53.43 48.31
MLDG [shu2019meta] 67.29 75.04 65.97 66.98 66.41 62.00 55.44 54.65 47.35 33.83 15.82 23.25 55.89 52.98 49.64
SWAD [cha2021swad] 68.58 78.86 68.33 63.29 65.39 58.28 51.68 52.20 44.65 26.48 24.19 16.84 52.51 55.16 47.03
MixStyle [zhou2020domain] 54.04 53.94 46.23 52.72 46.49 39.10 37.65 29.78 22.47 20.38 2.35 13.29 41.20 33.14 30.27
MEDIC-cls [wang2023generalizable] 62.76 69.29 60.03 63.10 62.25 55.06 56.42 56.30 49.08 32.16 8.13 24.16 53.61 48.99 47.08
MEDIC-bcls [wang2023generalizable] 62.76 50.05 44.19 63.10 45.67 46.56 56.42 39.96 40.02 32.16 27.74 23.91 53.61 40.86 38.67
EBiL-HaDS-cls [peng2024advancing] 62.84 64.75 57.55 63.79 62.07 56.86 48.07 45.94 38.18 36.00 10.31 24.98 52.68 45.77 44.39
EBiL-HaDS-bcls [peng2024advancing] 62.84 45.39 40.81 63.79 56.27 51.76 48.07 43.14 34.07 36.00 10.83 16.06 52.68 38.91 35.68
@@
Table 7: Results (%) of PACS on ViT-Base [dosovitskiy2021an]. The open-set ratio is 6:16{:}1 and symmetric label noise with a ratio of $80\%$ is selected.